OpenStudy (anonymous):

What kind of reason is used here? If a=b and c=d, then a+c=b+d. A. Deductive B. Inductive C. Hypothesis D. Conclusion

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

but wait** haha

OpenStudy (anonymous):

Oh okay haha thanks

OpenStudy (anonymous):

I thought it says inductive reasoning, then I'd be correct :) you're welcomed.

OpenStudy (anonymous):

no, the type of reasoning here is deductive. Inductive reason can be thought of as using a small subset of examples to prove a rule. Here you are given two rules and deducing the conclusion (which isn't a type of reasoning), thus are using deductive reasoning.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

hypothesis and conclusion are not types of reasoning, they are part of the method to create an experiment (or a mathematical). Those are trick answers.

OpenStudy (anonymous):

Oh okay thank you! Do you think you could answer this other one I have?@sylbot

OpenStudy (anonymous):

@sylbot I think he's right : o deductive reasoning is using a one step process. If... then..

OpenStudy (anonymous):

good job catching that :))

OpenStudy (anonymous):

yeah, shoot I should be able to help

OpenStudy (anonymous):

What kind of reasoning is Althea using? Althea notices that 4=2+2,6=3+3,8=5+3, and 28=23+5, so she thinks it is likely that she can write every number as the sum of two prime numbers. A. Deductive B. Inductive C. Hypothesis D. Conclusion

OpenStudy (anonymous):

This is a perfect example of what I was talking about earlier when I said inductive reasoning. Here she is using a small subset of examples to prove a rule. Each equation is an example, and the rule is " she can write every number as the sum of two prime numbers. "

OpenStudy (anonymous):

Oh okay so it's inductive reasoning

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Okay what about this

OpenStudy (princeharryyy):

i think u study alot hehe :)

OpenStudy (anonymous):

Refer to the following conditional statement: if a and b are odd integers then a-b is an even integer State the contrapositive of the statement. 1. If a- b is an even integer, then a and b are not odd integers 2. If a and b are not odd integers then a-b is not an even integer 3. If a-b is an even integer then a and b are odd integers 4. a-b is an even integer

OpenStudy (anonymous):

ok so do you know the general case of contrapositive? Essentially is the complete opposite of the positive. So the positive looks like If a, then b. The contrapositive looks like If not b, then not a. Can you figure it out with this?

OpenStudy (anonymous):

I think it's 3

OpenStudy (anonymous):

Oh nvm Idk lol

OpenStudy (anonymous):

nope thats the negative and its a very dangerous trap. Its not necessarily true.

OpenStudy (anonymous):

You just deleted it lol

OpenStudy (anonymous):

yeah, uh none of there choices are correct. The statement should be something like: "If a-b is not an integer, then a and b are not odd integers." This is the true contrapositive of the statement. I'm pretty certain that none of these choices are right.

OpenStudy (anonymous):

I deleted it because I needed to check myself real quick :P but yeah, http://www.mathwords.com/c/contrapositive.htm It seems like what I said holds true.

OpenStudy (anonymous):

Also when I said 3 was the negative, I mean 3 was the INVERSE, sorry about that. The inverse occurs when negating both sides of the logic statement but doesn't not necessarily hold true.

OpenStudy (anonymous):

Oh we'll then great lol

OpenStudy (anonymous):

Okay. No worries thank you :)

OpenStudy (anonymous):

did you get the answer? Was there a mistake or something?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Oh quick edit, I missed a word. The contrapositive should be "If a-b is not an even integer, then a and b are not odd integers"

OpenStudy (anonymous):

Yeah I saw that I just figured that is what you meant. :) lol